OpenStudy (anonymous):

Find the vertex: y = −2x^2 + 4x − 3 I already know that the answer is (1, -1) but I can't figure out how to solve it. This is what I did: y=ax^2+bx+c so y = −2x^2 + 4x − 3 is already in standard form. a=2, b=4, and c=3 x = -b/a = -4/2 = -2 (By now, I already know I must've done something wrong.) Substitute x back into the equation: y = −2(-2)^2 + 4(-2) − 3 (2)4 -8 - 3 8 -8 - 3 y= -3 So that means my vertex would be (-2, -3) but I know that it's wrong. Could someone help me figure out how to solve it?

OpenStudy (raden):

use x = -b/2a not -b/a

OpenStudy (raden):

-b/a used to find the sum of roots that equation : x1+x2 = -b/a

OpenStudy (harsimran_hs4):

a = -2, b=4, and c=-3 not a =2, b=4, and c=3

OpenStudy (anonymous):

But that time I got -5 y = −2x^2 + 4x − 3 a = 2, b = 4, and c = -3 x = -b/2a -4/2 * 2 = -1 substitute x back into the equation: y = −2(-1)^2 + 4(-1) − 3 -2(-1) - 4 - 3 2 -7 y = -5

Parth (parthkohli):

(-1)^2 = 1, not -1

OpenStudy (raden):

x = -b/2a x= -4/2 * (-2) = -4/-4 = 1

OpenStudy (anonymous):

So then I don't have to substitute x back into the equation?

Parth (parthkohli):

You get \(1\) for the \(x\)-coordinate, now substitute \(1\) for \(x\) to get the \(y\)-coordinate.

OpenStudy (anonymous):

Oh okay, I have it now. Thanks!

Parth (parthkohli):

NP!

OpenStudy (agent0smith):

Be careful with the signs... from your earlier post: y = −2x^2 + 4x − 3 a = 2, b = 4, and c = -3 a = -2 there, not positive 2.

OpenStudy (anonymous):

Thanks for pointing that out, I will remember next time :)